OpenStudy (anonymous):

okay one more confusion. how does 3^(n+1) divided by 3^(n+2) simplify to just plain old 3???

OpenStudy (anonymous):

it doesn't it simplifies to 1/3

OpenStudy (anonymous):

oh. so I guess I had it wrong. how would I go about simplifying the expression?

OpenStudy (anonymous):

\[3^{n+1}/3^{n+2} = 3^n \times 3^1 \div s^n 3^2 = 1/3\]

OpenStudy (anonymous):

remember ur rules of adding and subtracting exponents

OpenStudy (anonymous):

3^(n+2) = 3^(n+1+1) = 3^(n+1) * 3 so that why 3^(n+1) divided by 3^(n+2) = 3^(n+1) / 3^(n+1) * 3 =1/3

OpenStudy (anonymous):

so it equals 1/3. NOT just 3?

OpenStudy (anonymous):

yes it equals 1/3

OpenStudy (anonymous):

yep. i guess that's how it works.

OpenStudy (anonymous):

okay I am going to look over it and try to understand it. thank you both for the thorough explanations! :) much appreciated.

OpenStudy (anonymous):

welcome ;)